OpenStudy (anonymous):

Find the distance between the coordinates: 1. (2,1) & (8,9) 2. (5,6) & (0,3) 3. (-5,4) & (7,16)

OpenStudy (anonymous):

for each one use \[\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] for example the first one is \[\sqrt({8-2)^2+(9-1)^2}=\sqrt{6^2+8^2}=\sqrt{36+64}=\sqrt{100}=10\]

OpenStudy (anonymous):

So number one is 10?